OpenStudy (fanduekisses):

CAN SOMEONE HELP ME WITH THESE PHYSICS QUESTIONS? I'm reviewing for an exam...

OpenStudy (fanduekisses):

A net force of 10 N causes a mass to accelerate at 6 m/s2. What would be the acceleration if the mass were doubled and the force kept constant?

OpenStudy (acxbox22):

f=ma 10=6m 5/3=m 5/3 x 2= 10/3 10=10/3a a=??

OpenStudy (acxbox22):

do you understand the steps?

OpenStudy (mathmath333):

\(\large \tt \begin{align} \color{black}{F=ma\\~\\ 10=m6\\~\\ m=\dfrac{5}{3}\\~\\ \normalsize \text{now the mass is doubled but force kept constant}\\~\\ F=2ma\\~\\ 10=2\times \dfrac{5}{3}a\\~\\ }\end{align}\)

OpenStudy (acxbox22):

both ways represent the same thing @Fanduekisses

OpenStudy (fanduekisses):

it's 3 m.s :)

OpenStudy (fanduekisses):

3 m/s

OpenStudy (acxbox22):

yep 3m/s^2

OpenStudy (acxbox22):

;)

OpenStudy (fanduekisses):

hehe I forgot about that force is mass x acceleration

OpenStudy (acxbox22):

the acceleration was halved

OpenStudy (fanduekisses):

A 2,000-kilogram automobile accelerates horizontally to a speed of 20 m/s. The increase in its gravitational potential energy is:

OpenStudy (fanduekisses):

wait, wouldn;t it be zero? because it is moving...

OpenStudy (acxbox22):

post the questions separately so you can receive maximum help from more people, thanks ;)